Your grandfather would like to share some of his fortune with you. He offers to give you money under one of the following scenarios (you get to choose): Use PV table for $1 and Present Value of Ordinary Annuity of​ $1 table

1. $8,000 per year at the end of each of the next eight years

2.   $49,950 (lump sum) now

3. $100,050 (lump sum) eight years from now

Requirements:

1.

Calculate the present value of each scenario using an 8 % discount rate. Which scenario yields the highest present value? Round to the nearest whole dollar.

2.

Would your preference change if you used a 10 % discount rate?
Answer #1

Calculate present value :

1) Present value of option 1 = 8000*5.74664 = $45973.12 or $45973

2) Present value of option 2 = 49950*1 = 49950

3) Present value of option 3 = 100050*0.54027 = $54054

Option 3 has highest present value

Calculate following on 10%

1) Present value of option 1 = 8000*5.33493 = $42679

2) Present value of option 2 = $49950

3) Present value of option 3 = 100050*0.46651 = $46674

Yes in 10% option 2 has highest present value